OpenStudy (anonymous):

design a process of teaching using constructivist approach for the following problems: 2 gift boxes are with ribbon( cube and cylinder). Box a is a cube of side-length 10cm.Box b is a cylinder with height end diameter 10cm each. How would you teach your students to answer the question which box needs the longer ribbon?

OpenStudy (anonymous):

the box having more surface area will require longer ribbon. so we have to find out the surface area of both boxes. surface area of a cube is 6*l^2 (where l=side lenght) =6*10*10 =600 surface area of a cylinder is (area of bases +area of curved surface) = 2*(pi)*r^2+2*(pi)*r*l =2*3.14*10*10 +2*3.14*10*10 =1256 so the cyllinder has more surface area & need longer ribbon.

OpenStudy (anonymous):

tq for ur answer. bt hwto teach students abt it?